Message: 5) A car accelerates at 4 meters/s/s. Assuming the car starts from rest, how far will it travel in 10
s?
i use equation of kinamtic formula and it is wrong why?
D=vit+.5at^2 ?
why it is not correct to use this formula?
vi= 0
1
Expert's answer
2020-02-18T09:39:42-0500

Given:

"v_i=0\:\rm m/s;"

"a=4\:\rm m/s^2;"

"t=10\:\rm s."

The distance


"d=v_it+\frac{at^2}{2}=0\times 10+\frac{4\times 10^2}{2}=200\:\rm m."